# Unsupervised Elicitation of Language Models

### v1

@@ Abstract
To steer pretrained language models for downstream tasks, today’s post-training paradigm relies on humans to specify desired behaviors. However, for models with superhuman capabilities, it is difficult or impossible to get high-quality human supervision. To address this challenge, we introduce a new unsupervised algorithm, Internal Coherence Maximization (ICM), to fine-tune pretrained language models on their own generated labels, without external supervision. On GSM8k-verification, TruthfulQA, and Alpaca reward modeling tasks, our method matches the performance of training on golden supervision and outperforms training on crowdsourced human supervision. On tasks where LMs’ capabilities are strongly superhuman, our method can elicit those capabilities significantly better than training on human labels. Finally, we show that our method can improve the training of frontier LMs: we use our method to train an unsupervised reward model and use reinforcement learning to train a Claude 3.5 Haiku-based assistant. Both the reward model and the assistant outperform their human-supervised counterparts.
Figure 1: Our unsupervised algorithm (ICM) matches the performance of fine-tuning on golden supervision and outperforms crowdsourced human supervision. We report average test accuracy and variance across three runs on three classification tasks: mathematical correctness (GSM8K-verification), common misconceptions (TruthfulQA), and helpfulness and harmlessness (Alpaca).
@@ 1 Introduction
Today’s post-training paradigm of pre-trained language models (LMs) still relies on humans to specify desired behaviors, either through demonstrations or preference feedback [24, 14, 4]. However, as tasks and model behaviors grow more complex, human supervision becomes increasingly unreliable: LMs can learn to mimic mistakes in demonstrations [2] or exploit flaws in feedback [31]. How do we train LMs to do tasks that are too difficult for humans to demonstrate or evaluate reliably?
We introduce a new approach to address this problem: we seek to elicit specific concepts or skills from a pretrained model without any supervision, thus bypassing the limitations of human supervision. Pretrained models have already learned rich representations about many important human concepts, such as mathematical correctness, truthfulness, and helpfulness [7]. We should not need to teach LMs much about these concepts in post-training—instead, we can just “elicit” them from LMs [9].
@@
Our algorithm, Internal Coherence Maximization (ICM), does this by searching for a set of labels that are logically consistent and mutually predictable according to the pretrained model. Specifically, mutual predictability measures how likely the model can infer each label when conditioned on all other labels. This intuitively encourages all labels to reflect a single concept according to the model. Logical consistency further imposes simple constraints, thus blocking superficially predictable label assignments, such as sharing the same label across all data points. Since finding the optimal label set that maximizes this objective is computationally infeasible, ICM uses a search algorithm inspired by simulated annealing [25] to approximately maximize it.
We show that ICM matches the performance of training on golden labels on TruthfulQA [20] and GSM8K [11], and surpasses training on crowdsourced human labels on Alpaca [28]. Additionally, on a task where LMs are strongly superhuman—identifying an author’s gender from a writing sample [^1] We use a widely-adopted academic dataset [27] for studying AI fairness [10, 21], which consists of self-reported author information. ICM significantly outperforms the human supervision baseline.
Beyond standard benchmarks, we investigate ICM’s potential in improving frontier models by training a version of Claude 3.5 Haiku without any human supervision. Specifically, we first use ICM to train an unsupervised reward model (RM), then fine-tune the Claude 3.5 Haiku pretrained model through reinforcement learning. Evaluations on Rewardbench [18] confirm that our unsupervised RM outperforms its counterparts trained on production-grade high-quality human supervision. Further, when assessed by Claude 3.5 Sonnet’s production-grade RM, our unsupervised assistant policy wins 60% of head-to-head comparisons against the policy trained with the human-supervised RM.
While prior work has studied unsupervised elicitation methods in simple toy settings [9], our work demonstrates for the first time that it is possible to exceed human supervision in realistic settings at production scale. By successfully training a Claude 3.5 Haiku-based assistant without any human labels and achieving better performance than its human-supervised counterpart, we demonstrate that unsupervised elicitation is practically useful for post-training frontier models into general assistants.
@@ 2.2 Scoring Function
Mutual Predictability. For each example $x_{i}$, we calculate the probability of its label $y_{i}$ by putting all other N1N-1 labels in the context, and sum the log probabilities across all examples:
@@
where $P_{\theta}$ is the pretrained model.
Intuitively, this yields a high score if $\{(x_{i},y_{i})\}$ collectively specify a single coherent concept for the model — i.e. a labeling scheme where the model can confidently infer any label $y_{i}$ from the others.
However, mutual predictability alone allows some degenerate solutions due to artifacts of in-context learning, e.g. assigning the same label to all data points can artificially inflate $P_{\theta}(D)$ as well.
Logical Consistency. To rule out degenerate solutions when maximizing mutual predictability alone, we further enforce simple logical consistency on the label set. Specifically, we are given a logical consistency function $c(x_{i},y_{i},x_{j},y_{j})\in\{0,1\}$ that checks whether the labels $y_{i}$ and $y_{j}$ on data points $x_{i}$ and $x_{j}$ are logically consistent with each other. We use it to measure inconsistencies in our labels:
$\mathcal{I}(D)=\sum_{i=1}^{N}\sum_{j=1}^{N}c(x_{i},y_{i},x_{j},y_{j})$
Determining fine-grained logical consistency between each example is non-trivial; however, empirical evidence suggests that even simple and general logical constraints suffice. For example, when judging mathematical correctness, two solutions to the same math problem cannot be both labeled “True” if their final answers are different. Another general, task-agnostic logical constraint that we use for comparative datasets is asymmetry: when comparing two responses $A$ and $B$, two claims “A>BA>B” and “B>AB>A” cannot be both labeled “True”.
@@ 2.3 Our Algorithm
Finding the optimal label set that maximizes our scoring function is an integer programming problem, which is computationally infeasible for realistic dataset sizes ($10^{3}<N<10^{6}$). ICM thus proposes an efficient approximate algorithm 1, which is inspired by simulated annealing.
Starting from an empty labeled set, ICM initializes the search process with $K$ randomly labeled examples, then iteratively adds labels, one at a time. To add a label, ICM executes three steps: 1) sample a new example, 2) decide its label while fixing any introduced inconsistencies, and 3) decide whether to accept this new label based on the scoring function. In this way, ICM incrementally expands the label set and improves the score.
@@
Initialization. We initialize the searching process with $K$ randomly labeled examples. The choice of $K$ presents a trade-off. A large $K$ (e.g., $K=N$) introduces significant initial noise that hinders subsequent convergence. Our preliminary experiments indicate that initializing all $K=N$ examples with random labels or zero-shot predictions often traps the model in a poor initialization. Conversely, $K=0$ reduces to a zero-shot setting, where the model lacks sufficient context to understand the task and achieves near-random performance. Empirically, we find that a small number (e.g., $K=8$), often strikes a good balance by providing sufficient demonstrations while reducing initial noise [23].
@@ 3.1 Datasets
See Table 1 for dataset examples. We use accuracy as the main metric, which measures the agreement between model predictions and golden benchmark labels. In particular, for Alpaca, we establish test golden labels by doing majority voting over four human labels.
@@ 3.2 Baselines
We adopt the following four baselines in our experiments:
  • Zero-shot indicates zero-shot prompting on pretrained models. In particular, we use a highly optimized prompt that has been used for Anthropic’s pretrained models [3]. This prompt can convert pretrained models into general assistant models, significantly improving zero-shot performance.
  • Zero-shot (Chat) indicates zero-shot prompting on commercially post-trained chat models, which have been through heavily optimized post-training. As an example, the llama-2 chat models are post-trained on nearly 30K human demonstrations and 3 million human preference feedback [29].
  • Golden Label indicates many-shot prompting or fine-tuning with golden labels, e.g., labels from TruthfulQA and GSM8K.
@@
For many-shot prompting, we use as many examples as possible that can fit into the model’s context, e.g., 160 examples for Alpaca.
3.3 Models
In our experiments we use two open-weight models, Llama 3.1 8B and Llama 3.1 70B, and two proprietary models, Claude 3 Haiku and Claude 3.5 Haiku. Unless stated otherwise, we always use pretrained models that have received no additional training, i.e. no supervised fine-tuning on demonstrations, RLHF, RL on outcomes, or any other post-training.
@@ 4.1 Eliciting Capabilities on Common NLP Tasks
Figure 3: Results with Llama 3 pretrained models, 8B for GSM8K, 70B for TruthfulQA and Alpaca.
@@ 4.2 Unsupervised Elicitation Fails when Concepts are not Salient
To highlight some of our algorithm’s limitations, we design a task specifically to be impossible for unsupervised elicitation. Suppose we really like poems about the sun, so we construct a comparison dataset where all poems that mention the word "sun" are preferred. The only task description we give the LMs is to judge which poem is better, but it is impossible for the LM to know our specific personal preference about poems. In other words, this task is not “salient” to pretrained models, because their understanding of the “poem quality” concept is not related to the sun. To construct the dataset, we use Claude 3.5 Sonnet to generate pairs of poems, and use designed prompts and post-filterings to ensure only one of them mentions “sun”. Experiment results with Llama 70B are shown in Figure 5. As expected, we find ICM performs no better than random guessing.
@@ 4.3 Eliciting Superhuman Capabilities
After studying unsupervised elicitation on three common NLP datasets, we are further interested in tasks where pretrained models are strongly superhuman. To study this, we explore an author gender prediction task using the Blog Authorship Corpus [27]. [^2] Our goal is not to improve AI performance at predicting author gender, but rather to study how well this capability is already present in pretrained models.
@@ 4.4 Training an Assistant Chatbot without Supervision
Figure 7: Accuracy of reward models (left) and pairwise winrates of assistant policy models against the human-supervised baseline (right). We train a Claude 3 Haiku-based reward model, using the Alpaca data or the production data used for training publicly released Claude 3.5 Haiku. Next, we optimize the Claude 3.5 Haiku pretrained model against our reward model to build an assistant policy.
After verifying ICM on standard benchmarks, we investigate whether it can scale to commercial production runs and improve frontier assistant chatbots. Specifically, we aim to train a helpful chat assistant based on the Claude 3.5 Haiku pretrained model, without introducing any human preferences or supervision labels whatsoever.
Reward Model Training. We use Claude 3 Haiku [^3] This was an oversight on our part. Ideally, we would use the same model for reward model and policy. to generate unsupervised labels. We use the task description “which response is more helpful, harmless, and honest?”. We sample a subset from the production preference dataset for training the publicly released Claude 3.5 Haiku. This subset consists of nearly 400K examples with a 64K token limit. We first use ICM to label 6K examples, train an initial reward model (RM) to label the rest of the data, and then train the final unsupervised RM. We also run the same process on Alpaca to serve as a baseline against this production data.
We conduct evaluations on Rewardbench [18], a widely-used challenging benchmark for RMs. Figure 7 (left) shows the results. First, the human-supervised RM trained on the production data significantly outperforms the RM trained on Alpaca, due to its high-quality human labels and complex examples. Consequently, surpassing the human-supervised RM trained on production data is much harder. Nevertheless, our unsupervised RM still achieves higher accuracy (75.0% v.s. 72.2%).
Reinforcement Learning with Unsupervised RM. Using both the unsupervised and human-supervised RM, we train two policies via reinforcement learning to create helpful, harmless, and honest assistants. We train both policies on 20,000 RL episodes. We conduct head-to-head comparisons between two policies: each model’s responses are graded by the RM for training the publicly released Claude 3.5 Sonnet model. As shown in Figure 7 (right), the policy trained with the unsupervised RM achieves a 60% win rate. Both these policies lag severely behind the performance of the publicly released Claude 3.5 Haiku, which achieves a much higher 92% win rate against the human-supervised baseline. This is expected because the publicly released Claude 3.5 Haiku is trained for much longer on a much larger dataset with a Claude 3.5 Haiku-based RM. Overall, these experiments suggest that ICM can scale to commercial production runs.
@@ 5 Ablations
Evaluating robustness to worst-case initialization. It is possible that our algorithm could collapse under bad initialization (e.g., all initial $K$ labels are wrong), but we coincidentally never encounter such scenarios in Sec. 4 because they happen rarely. We thus investigate ICM’s robustness against different initializations:
  • Golden: using golden dataset labels. This corresponds to a semi-supervised setting.
  • Random: using random labels (our default setting).
  • Worst: using entirely wrong labels.
@@
Figure 9 showcases results on TruthfulQA with the Llama 8B model. We report the test accuracy using many-shot prompting. Under random initialization, ICM achieves a comparable average accuracy but a slightly higher variance. Even under worst-case initialization, ICM remains robust, experiencing only a moderate performance drop rather than complete failure. This is mainly due to its iterative nature: a few initial bad labels would not degrade the performance significantly, as they can be gradually corrected as the algorithm progresses.
@@
Ablating logical consistency. The logical consistency term may be of limited value: ICM only introduces simple and general logical consistency that can be applied to many tasks, because determining fine-grained consistency relationships across examples is challenging. Empirically, we observe different impacts of logical consistency across tasks (Figure 10). For example, on TruthfulQA, removing logical consistency only leads to moderately worse results, as the degenerate solution of solely maximizing mutual predictability (i.e. assigning the same label everywhere) happens rarely. In contrast, logical consistency is crucial on Alpaca, since the degenerate solution almost always happens without that.
6 Related Work
Scaling beyond Human Supervision. Recent work has shown diverse failure modes of post-training LMs with unreliable human supervision. For example, LMs can learn to reward-hack human-designed supervision signals [6] or even real humans themselves [31]. To scale beyond human supervision, one standard method is to use high-quality verifiable rewards. For example, in math, we can match model outputs with existing ground truth solutions [15]. Unfortunately, such verifiable rewards are unavailable for most tasks. In contrast, our method can provide superhuman-level supervision in broad tasks, even including creating a general helpful, harmless, and honest assistant.
Evidence of Latent Capabilities in LMs. Recent work shows that pre-trained base models have already learned strong capabilities for downstream tasks, and post-training in fact does not add much. For example, pretrained models can achieve a comparable or even higher pass@$k$ than their post-trained counterparts when $k$ is large enough, even when post-training is done with verifiable rewards [32]. Similarly, pretrained and post-trained models perform nearly identically in decoding, while most distribution shifts occur with stylistic tokens such as discourse markers [19]. When inspecting model latent representations, recent work also finds that LMs encode strong signals of reasoning correctness [33] or hallucination [17, 13]. However, despite prior empirical evidence about LMs’ latent capabilities, they still fail to elicit them effectively.
Unsupervised Elicitation of LMs. CCS [9] is one of the most representative works for unsupervised elicitation, which works by solely using simple logical consistency to find latent knowledge. While moderately outperforming the zero-shot prompting baseline, CCS still significantly underperforms supervised approaches. As argued in [12], CCS, as well as other unsupervised approaches, often cannot find knowledge, because there are many other prominent features that can satisfy logical consistency properties. Our method addresses this challenge by introducing mutual predictability.
Several concurrent studies explore unsupervised elicitation by minimizing label entropy [34, 1], differing from our scoring function. Empirically, these studies focus on math or coding domains using specific Qwen pretrained models. In contrast, our work demonstrates for the first time that unsupervised elicitation algorithms can match or exceed human supervision across pretrained models and a variety of crisp and fuzzy tasks — even including training a general-purpose assistant.
Unsupervised elicitation can also be thought of as a special case of weak-to-strong generalization [8, 16]: while they try to use weak human supervision to elicit strong LMs, we seek to ignore the weak human supervision altogether.

### v2

@@ Abstract
To steer pretrained language models for downstream tasks, today’s post-training paradigm relies on humans to specify desired behaviors. However, for models with superhuman capabilities, it is difficult or impossible to get high-quality human supervision. To address this challenge, we introduce a new unsupervised algorithm, Internal Coherence Maximization (ICM), to fine-tune pretrained language models on their own generated labels, without external supervision. On GSM8k-verification, TruthfulQA, and Alpaca reward modeling tasks, our method matches the performance of training on golden labels and outperforms training on crowdsourced human supervision. On tasks where LMs’ capabilities are strongly superhuman, our method can elicit those capabilities significantly better than training on human labels. Finally, we show that our method can improve the training of frontier LMs: we use our method to train an unsupervised reward model and use reinforcement learning to train a Claude 4 Sonnet-based assistant. The resulting assistant matches its counterpart trained on production-grade human labels on average, with higher scores on chat and safety yet lower scores on math and coding.
Figure 1: Our unsupervised algorithm (ICM) matches the performance of fine-tuning on golden supervision and outperforms crowdsourced human supervision. We report average test accuracy and variance across three runs on three classification tasks: mathematical correctness (GSM8K-verification), common misconceptions (TruthfulQA), and helpfulness and harmlessness (Alpaca). Results are based on Llama 3 pretrained models, 8B for GSM8K, 70B for TruthfulQA and Alpaca.
@@ 1 Introduction
Today’s post-training paradigm of pre-trained language models (LMs) still relies on humans to specify desired behaviors, either through demonstrations or preference feedback (Ouyang et al., 2022; Glaese et al., 2022; Bai et al., 2022a). However, as tasks and model behaviors grow more complex, human supervision becomes increasingly unreliable: LMs can learn to mimic mistakes in demonstrations (Asare et al., 2023) or exploit flaws in feedback (Wen et al., 2024b). How do we train LMs to do tasks that are too difficult for humans to demonstrate or evaluate reliably?
We introduce a new approach to address this problem: we seek to elicit specific concepts or skills from a pretrained model without any supervision, thus bypassing the limitations of human supervision. Pretrained models have already learned rich representations about important human concepts, such as mathematical correctness, truthfulness, and helpfulness (Burns et al., 2022). We should not need to teach LMs much about these concepts in post-training—instead, we can just “elicit” them from LMs.
@@
Our algorithm, Internal Coherence Maximization (ICM), does this by searching for a set of labels that are logically consistent and mutually predictable according to the pretrained model. Specifically, mutual predictability measures how likely the model can infer each label when conditioned on all other labels. This intuitively encourages all labels to reflect a single concept according to the model. Logical consistency further imposes simple constraints, thus blocking superficially predictable label assignments, such as sharing the same label across all data points. Since finding the optimal label set that maximizes this objective is computationally infeasible, ICM uses a search algorithm inspired by simulated annealing (Pirlot, 1996) to approximately maximize it.
We show that ICM matches the performance of training on golden labels on TruthfulQA (Lin et al., 2021) and GSM8K (Cobbe et al., 2021), and surpasses training on crowdsourced human labels on Alpaca (Taori et al., 2023). Additionally, on a task where LMs are strongly superhuman—identifying an author’s gender from a writing sample [^1] We use a widely-adopted academic dataset (Schler et al., 2006) for studying AI fairness (Coavoux et al., 2018; Lyu et al., 2020), which consists of self-reported author information.ICM significantly outperforms human supervision.
Beyond standard benchmarks, we investigate ICM’s potential in improving frontier models by training a version of Claude 4 Sonnet assistant without any human supervision. Specifically, we first train a reward model (RM) with ICM, then train an assistant via reinforcement learning, which is assessed by Claude 4 Opus’s production-grade RM. Compared with the counterpart trained on production-grade human labels, our unsupervised assistant learns faster during RL and yields comparable scores on average, with higher scores on chat and safety and lower scores on math and code.
While prior work has studied unsupervised elicitation methods in simple toy settings (Burns et al., 2022), our work demonstrates for the first time that it is possible to match or exceed human supervision in realistic settings. By successfully training a Claude 4 Sonnet-based assistant without any human labels and achieving comparable performance to its human-supervised counterpart, we show that unsupervised elicitation is practically useful for post-training frontier models into general assistants.
@@ 2.1 Problem Statement
In particular, we are mainly focused on classification tasks (e.g. reward modeling) in this paper, as we can naturally use reinforcement learning to optimize for open-ended generation tasks. We demonstrate this by training a version of genreal Claude 4 Sonnet assistant in Sec. 4.4.
@@ 2.2 Scoring Function
Mutual Predictability. As illustrated in the top panel of Figure 2, given a pretrained model $P_{\theta}$, for each example $x_{i}$, we calculate the probability of its label $y_{i}$ by conditioning all other |D|1|D|-1 labels (e.g. via in-context learning as in Algorithm 1), and sum the log probabilities across all examples:
@@
Intuitively, this yields a high score if $\{(x_{i},y_{i})\}$ collectively specify a single coherent concept according to $P_{\theta}$, i.e. a labeling scheme where $P_{\theta}$ can confidently infer any label $y_{i}$ from the others. Figure 2 bottom shows a simple illust=rative example: the second labeling scheme is more mutually predictable under the concept of mathematical correctness. See Appendix C for another example.
However, mutual predictability alone allows some degenerate solutions, e.g. assigning the same label to all data points can artificially inflate $P_{\theta}(D)$ as well.
Logical Consistency. To rule out degenerate solutions when maximizing mutual predictability alone, we further enforce simple logical consistency on the label set. Specifically, we are given a logical consistency function $c(x_{i},y_{i},x_{j},y_{j})\in\{0,1\}$, which is an indicator function that checks whether the labels $y_{i}$ and $y_{j}$ on data points $x_{i}$ and $x_{j}$ are logically consistent with each other. We use it to measure inconsistencies in our labels:
$\mathcal{I}(D)=\sum_{i=1}^{|D|}\sum_{j=1}^{|D|}c(x_{i},y_{i},x_{j},y_{j})$
Determining fine-grained logical consistency between each example is non-trivial; however, empirical evidence suggests that even simple and general logical constraints suffice. For example, when judging mathematical correctness, two different answers cannot both be True. Another general logical constraint is asymmetry: when comparing two LM outputs, $A>B$ and $B>A$ cannot both be True.
@@ 2.3 Our Algorithm
Finding the optimal label set that maximizes our scoring function is an integer programming problem, which is computationally infeasible for realistic dataset sizes ($|D|>10^{3}$). ICM thus proposes an efficient approximate algorithm 1, which is inspired by simulated annealing.
Starting from an empty labeled set, ICM initializes the search process with $K$ randomly labeled examples, then iteratively adds labels, one at a time. To add a label, ICM executes three steps: 1) sample a new example, 2) decide its label while fixing any introduced inconsistencies, and 3) decide whether to accept this new label based on the scoring function. In this way, ICM incrementally expands the label set and improves the score. The bottom of Figure 2 illustrates this iterative process.
@@
Initialization. We initialize the searching process with $K$ randomly labeled examples. The choice of $K$ presents a trade-off. A large $K$ introduces significant initial noise that hinders subsequent convergence. Preliminary results show that initializing all $K=|D|$ examples with random labels or zero-shot predictions often traps the model in a poor initialization. Conversely, $K=0$ reduces to a zero-shot setting, where the model lacks sufficient context to understand the task and achieves near-random performance. Empirically, we find that a small number (e.g., $K=8$) often strikes a good balance by providing sufficient demonstrations while reducing initial noise (Min et al., 2022).
@@ 3.1 Datasets
See Table 1 for dataset examples. Regarding logical consistency checks, for GSM8K and TruthfulQA, we use “two different answers cannot both be true”. For Alpaca, we use ”$A>B$ contradicts B>AB>A”. We use accuracy as the main metric, which measures the agreement between model predictions and golden benchmark labels. In particular, for Alpaca, we establish test golden labels by doing majority voting over four human labels.
@@ 3.2 Baselines
We adopt the following four baselines in our main experiments. Appendix E compares ICM with more baselines (e.g. distilling from GPT-4o), and ICM consistently yields better performance.
  • Zero-shot indicates zero-shot prompting on pretrained models. In particular, we use a highly optimized prompt that has been used for Anthropic’s pretrained models (Askell et al., 2021). which converts pretrained models into general assistants, significantly improving zero-shot performance.
  • Zero-shot (Chat) indicates zero-shot prompting on commercial chat models, which have been through heavily optimized post-training. For example, Llama 2 chat models are post-trained on nearly 30K human demonstrations and 3 million human preference feedback (Touvron et al., 2023).
  • Golden Label indicates many-shot prompting or fine-tuning with golden labels. We use as many examples as possible that can fit into the model’s context, e.g., 160 examples for Alpaca.
@@ 4.1 Eliciting Capabilities on Common NLP Tasks
Figure 3: Prompting or fine-tuning results with Llama 3 models, 8B for GSM8K, 70B for the others.
@@ 4.2 Unsupervised Elicitation Fails when Concepts are not Salient
To highlight some of our algorithm’s limitations, we design a task specifically to be impossible for unsupervised elicitation. Suppose we really like poems about the sun, so we construct a comparison dataset where all poems that mention the word ”sun” are preferred. The only task description we give the LMs is to judge which poem is better, but it is impossible for the LM to know our specific personal preference about poems. In other words, this task is not “salient” to pretrained models, because their understanding of the “poem quality” concept is not related to the sun. To construct the dataset, we use Claude 3.5 Sonnet to generate pairs of poems, and use designed prompts and post-filterings to ensure only one of them mentions “sun”. Experiment results with Llama 70B are shown in Figure 5. As expected, we find ICM performs no better than random guessing.
@@ 4.3 Eliciting Superhuman Capabilities
After studying unsupervised elicitation on three common NLP datasets, we are further interested in tasks where pretrained models are strongly superhuman. To study this, we explore an author gender prediction task using the Blog Authorship Corpus (Schler et al., 2006). [^2] Our goal is not to improve AI performance at predicting author gender, but rather to study how well this capability is already present in pretrained models.
@@ 4.4 Training an Assistant Chatbot without Supervision
After verifying ICM on standard benchmarks, we investigate whether it can scale to commercial production runs and improve frontier assistant chatbots. Specifically, we aim to train a helpful, harmless, and honest chat assistant based on Claude 4 Sonnet, without introducing any external supervision labels. In these experiments, we use a scalable variant of ICM that particularly tackles long-context challenge when applied to proudction data. See Appendix G for more details.
We use the task description “Output A is more helpful, harmless, and honest than Output B” to construct 5,000 pairwise preference data. Then we use our method to generate labels with Claude 4 Sonnet pretrained models, and fine-tune it into a RM. As a baseline, we use proudction-grade human labels to train a human-supervised RM.
Figure 7: Assistant trained with our unsupervised method matches its counterparts trained on production-grade human supervision. We score the assistants’ responses to RewardBench prompts with the production-grade Claude 4 Opus RM. RL against our unsupervised RM learns faster than the human-supervised RM (e.g., 2.5x the speed on Chat and Chat-Hard).
Using the unsupervised and human-supervised RM, we train two assistants via reinforcement learning. The training data is a mix of math, code and instruction following tasks. We train both assistant policies on 250,000 episodes. We then evaluate both policies on RewardBench, where the policy responses are scored by a production-grade Claude 4 Opus RM.
Results are shown in Figure 7. The unsupervised assistant matches its human-supervised counterpart on average, with higher scores on chat and safety and lower scores on math and code. We suspect that this is because the production-grade human labels are of higher quality on these crisp reasoning tasks.Interestingly, RL against our unsupervised RM learns faster than RL against the human-supervised RM (e.g., 2.5x the speed on Chat and Chat-Hard).
@@ 5 Ablations
Evaluating robustness to worst-case initialization. It is possible that ICM could collapse under bad initialization (e.g., all initial $K$ labels are wrong), but we coincidentally never encounter that in Sec. 4 because it happens rarely.
@@
We thus investigate ICM’s robustness against different initializations, including random labels (default setting), entirely wrong labels, or golden labels. Figure 9 showcases results on TruthfulQA with the Llama 8B model. We report the test accuracy using many-shot prompting. Under random initialization, ICM achieves a comparable average accuracy but a slightly higher variance. Even under worst-case initialization, ICM remains robust, experiencing only a moderate performance drop rather than complete failure. This is mainly due to its iterative nature: a few initial bad labels would not degrade the performance significantly, as they can be gradually corrected as the algorithm progresses.
@@
Ablating logical consistency. Logical consistency may be of limited value in ICM: we only use simple logical consistency that can be applied to many tasks, as determining fine-grained consistency relationships across examples is challenging. Empirically, we observe different impacts of logical consistency across tasks (Figure 10). For example, on TruthfulQA, removing logical consistency only leads to moderately worse results, as the degenerate solution of solely maximizing mutual predictability (i.e. assigning the same label everywhere) happens rarely. In contrast, logical consistency is crucial on Alpaca, since the degenerate solution almost always happens without that.